OpenStudy (anonymous):

Miami Beach, Florida, has 86,916 residents. Of those, about 13 out of every 100 are under 18 years old. Approximately how many people in Miami Beach are under 18?

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (anonymous):

\[\frac{13}{100}=\frac{x}{86,916}\] \[.13=\frac{x}{86,916}\] \[x=86,916\times .13\] and a calculator